OpenStudy (jacobciezki):

A container holds 21 spoons and 27 forks. What is the ratio of spoons to the total number of utensils in the container? A. 7 to 9 B. 7 to 16 C. 9 to 7 D. 16 to 9

OpenStudy (jacobciezki):

help

OpenStudy (jacobciezki):

how

OpenStudy (anonymous):

The ratio it slooking for how many utensils are spoons. It is \[\frac{ number of spoons }{ number of spoons and forks }\] So, you would get 21/48, which simplifies to 7/16

OpenStudy (skullpatrol):

21/48

OpenStudy (anonymous):

Whoops, misread it. The fraction you need to simplify would be \[\frac{ 21 }{ 48 }\]

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @LastTrainHome22 The ratio it slooking for how many utensils are spoons. It is \[\frac{ number of spoons }{ number of spoons and forks }\] So, you would get 21/48, which simplifies to 7/16 \(\color{blue}{\text{End of Quote}}\) Let's add some spaces to avoid confusion: :)\[\frac{number~of~spoons}{number~of~spoons~and~forks}\]

OpenStudy (jacobciezki):

ther is no 21 to 48

OpenStudy (skullpatrol):

Simplify

OpenStudy (anonymous):

That is because you need to simplify the fractions first.

OpenStudy (anonymous):

Read @LastTrainHome22 full answer and you will see the answer. :)

OpenStudy (jacobciezki):

ok

OpenStudy (jacobciezki):

A fruit vendor has 28 pears and 16 oranges. What is the ratio of oranges to pears in simplest form?

OpenStudy (skullpatrol):

a to b = a/b

OpenStudy (anonymous):

\[\frac{ 16 }{ 28 } \tt{~Simplify!}\]

OpenStudy (jacobciezki):

i got o.57

OpenStudy (anonymous):

An easy way to simplify fractions of to keep dividing the top and bottom by 2 until you can't any more. So here you get \[\frac{ 16 }{ 28 } = \frac{ 8 }{ 14 } = \frac{ 4 }{ 7 }\]

OpenStudy (anonymous):

The answer should be a fraction.

OpenStudy (jacobciezki):

ok

OpenStudy (skullpatrol):

4 to 7 = 4/7

OpenStudy (jacobciezki):

its not 7/4

OpenStudy (skullpatrol):

Oranges to pears

OpenStudy (jacobciezki):

yes

OpenStudy (jacobciezki):

so its not 4/7

OpenStudy (skullpatrol):

It is 4/7 or 4 oranges to 7 pears

OpenStudy (jacobciezki):

i think it2/3

OpenStudy (jacobciezki):

right

OpenStudy (skullpatrol):

Check your work please :-)

OpenStudy (jacobciezki):

you cheak yours and how you do it

OpenStudy (jacobciezki):

A bake sale has 12 cookies, 9 pies, 8 cakes, 4 tarts, and 3 cobblers to sell. What is the ratio of the total number of baked goods to cookies in simplest form?

OpenStudy (skullpatrol):

(12 + 9 + 8 + 4 + 3)/12

OpenStudy (skullpatrol):

36/12 = 3/1

OpenStudy (skullpatrol):

Not 3 because it isn't a ratio

OpenStudy (jacobciezki):

yea

OpenStudy (skullpatrol):

Note that 3 is the simplest form of the fraction 3/1.

OpenStudy (jacobciezki):

i got 60%

OpenStudy (skullpatrol):

Keep trying, getting good at these takes a lot of practice.

OpenStudy (jacobciezki):

ye ill take it again

OpenStudy (jacobciezki):

can you help me again

OpenStudy (skullpatrol):

I'll try :-)

OpenStudy (jacobciezki):

ok thx

OpenStudy (skullpatrol):

Thanks for trying to learn.

OpenStudy (jacobciezki):

ok one sec

OpenStudy (jacobciezki):

you ready

OpenStudy (skullpatrol):

:-)

OpenStudy (jacobciezki):

A jar contains 11 pennies and 8 dimes. What is the ratio of dimes to pennies? A. 19 to 11 B. 11 to 19 C. 8 to 11 D. 11 to 8

OpenStudy (jacobciezki):

c right

OpenStudy (jacobciezki):

@tinylittlehelper

OpenStudy (skullpatrol):

C is right :D

OpenStudy (jacobciezki):

ok ill tag you